OpenStudy (anonymous):

Find the lateral area of a cylinder with base area of 49 Π in2 and a height that is twice the radius of its base. 196 Π in2 686 Π2 in2 169 Π in2 196 Π2 in2

OpenStudy (anonymous):

whats the area of a circle?

OpenStudy (anonymous):

i mean the formula for it, do you know it?

OpenStudy (anonymous):

Formula for the cylinder correct?

OpenStudy (anonymous):

No, let's start with the base first which is a circle and then move to the more complex one which is the cylinder area, ok? Circle Area = pi*r^2 here we have 49*pi, so r^2 = 49, can you find r = ? hint: use square root.

OpenStudy (anonymous):

I have to do sqrt 49 right?

OpenStudy (anonymous):

yes, what will that give you for r?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

what does that mean for the height h of the cylinder? h = ?

OpenStudy (anonymous):

That h = 7

OpenStudy (anonymous):

but the question says "height that is twice the radius of its base." did you read the question and understood what is required? this is the most important of solving the question - is to understand everything about it :)

OpenStudy (anonymous):

When it means twice we're talking about 7 x 7

OpenStudy (anonymous):

no dear twice means 7+7

OpenStudy (anonymous):

or 2 times 7

OpenStudy (anonymous):

you have $10 and I have twice what you have, how much do I have? 100 or 20?

OpenStudy (anonymous):

20$

OpenStudy (anonymous):

that's the meaning of twice

OpenStudy (anonymous):

now you know r = 7 and h = 14 find the circumference of the circle (or the base) using the formula 2*pi*r , can you work it out please?

OpenStudy (anonymous):

14 (pi)

OpenStudy (anonymous):

well done now the lateral area of a cylinder if you google it http://www.thefreemathtutor.com/lateralareaofacylinder.html is 2*pi*r*h, we know already what 2*pi*r is and we know h, so can you work it out please?

OpenStudy (anonymous):

196 (pi)

OpenStudy (anonymous):

196 (pi) in^2

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

How about finding the surface area?

OpenStudy (anonymous):

that will be \[2\pi*r ^{2} + 2\pi*rh\]

OpenStudy (anonymous):

294 (pi)^2

OpenStudy (anonymous):

medal for you

OpenStudy (anonymous):

How do I give a medal?

OpenStudy (anonymous):

press the best response button on the correct answer, u get to do it only once per question

OpenStudy (anonymous):

I have several questions to ask.

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

you'll have to post each separately and let everyone get a fair go to help :)

OpenStudy (anonymous):

I always bump it.